# Iterative Prompt Relocation for Distribution-Adaptive Visual Prompt Tuning

### v1

@@ Appendix H Visualization and Analysis
Figure ix: Visualization of attention maps. PRO-VPT exhibits more focused and precise attention with fewer artifacts.
Learned Distributions and Accuracy Curves. Fig. viii illustrates the learned distributions in PRO-VPT as well as the accuracy curves in comparison to VPT, based on 100 training epochs. Empirical results from more datasets further reinforce that the prompting importance for each block is inherently task-dependent. Moreover, the comparison of accuracy curves between PRO-VPT and VPT, particularly on the SVHN, Camelyon, and Resisc45 datasets, reveals that PRO-VPT still exhibits an upward trend in accuracy during the late training stages. This highlights the effectiveness of learning the optimal distribution for visual prompts, which unlocks their full potential and maximizes task performance.
@@
Attention Maps. We also visualize the attention maps between [CLS] and image patches on VTAB-1k Natural Cifar100. As shown in Figs. ix(a)-(c), while VPT successfully focuses on the object, its attention exhibits significant artifacts [4] and, more critically, remains scattered. For example, in Fig. ix(a), VPT shows certain attention on the lake, which is actually part of the background. In contrast, PRO-VPT demonstrates more focused and accurate attention with fewer artifacts. Furthermore, as illustrated in Figs. ix(d) and (e), VPT appears to struggle with effectively concentrating on the object, whereas PRO-VPT maintains its ability to focus on the object.
Appendix I Limitation and Future Work
@@
Despite demonstrating promising performance and enhanced robustness, our approach still has certain limitations. First, as noted in [53, 19], prompt-based methods are significantly sensitive to the number of prompts per block. Although our method improves robustness and mitigates this issue to some extent (see § 4.5), there is still room for further refinement. In particular, while PRO-VPT maintains robust performance within a certain range (e.g., 10-50 initialized prompts per block as in Fig. vi), its performance still deteriorates when the number of prompts deviates significantly from the optimal value. Consequently, our method still employs the optimal number of prompts per task from [20], resulting in relatively high parameter counts akin to VPT’s. Second, the constraint of relocating one prompt per iteration (see § D) restricts our method to relocating only limited prompts within certain epochs. This particularly affects tasks requiring high prompt numbers, hindering PRO-VPT’s ability to learn optimal distributions. For instance, the changes from initial uniform to learned prompt distributions are less pronounced for Camelyon and DMLab as shown in Fig. viii, which require 100 initialized prompts per block, compared to other tasks. Nevertheless, it is worth noting that PRO-VPT still delivers considerable performance enhancements on these datasets. Addressing these limitations would further enhance the efficiency and applicability of our approach in diverse computer vision tasks.
@@
Moreover, while this paper focuses on calibrating the distribution of prompts, the PRO-VPT system is applicable to most block-wise PEFT approaches, e.g., adjusting the pre-block dimension of subspaces (i.e., the rank distribution) in adapter-based methods. An essential future direction deserving of further investigation is integrating our approach with other PEFT methods, developing a universal framework for fine-grained robust PEFT.

### v2

@@ Appendix I Visualization and Analysis
Figure viii: Visualization of attention maps. PRO-VPT exhibits more focused and precise attention with fewer artifacts.
Figure ix: Visualization of the prompt distributions learned by PRO-VPT and the accuracy curves compared to VPT on the VTAB-1k datasets: Natural DTD, SVHN; Specialized Camelyon, Resisc45; and Structured DMLab, KITTI-Dist.
@@
Attention Maps. We visualize the attention maps between [CLS] and image patches on VTAB-1k Natural Cifar100. As shown in Figs. viii(a)-(c), while VPT successfully focuses on the object, its attention exhibits significant artifacts [4] and, more critically, remains scattered. For example, in Fig. viii(a), VPT shows certain attention on the lake, which is actually part of the background. In contrast, PRO-VPT demonstrates more focused and accurate attention with fewer artifacts. Furthermore, as illustrated in Figs. viii(d) and (e), VPT appears to struggle with effectively concentrating on the object, whereas PRO-VPT maintains its ability to focus on the object.
Learned Distributions and Accuracy Curves. Fig. ix illustrates the learned distributions in PRO-VPT as well as the accuracy curves in comparison to VPT, based on 100 training epochs. Empirical results from more datasets further reinforce that the prompting importance for each block is inherently task-dependent. Moreover, the comparison of accuracy curves between PRO-VPT and VPT, particularly on the SVHN, Camelyon, and Resisc45 datasets, reveals that PRO-VPT still exhibits an upward trend in accuracy during the late training stages. This highlights the effectiveness of learning the optimal distribution for visual prompts, which unlocks their full potential and maximizes downstream performance.
Appendix J Limitation and Future Work
@@
Despite demonstrating promising performance and enhanced robustness, our approach still has certain limitations. First, as noted in [58, 21], prompt-based methods are significantly sensitive to the total number of prompts. Although our method improves robustness and mitigates this issue to some extent (refer to § 4.5), there is still room for further refinement. In particular, while PRO-VPT maintains robust performance within a certain range (e.g., 120-600 initialized prompts as in Fig. vi), its performance still deteriorates when the number of prompts deviates significantly from the optimal value. Consequently, our method still employs the optimal total number of prompts per task from [22], resulting in relatively high parameter counts as in VPT. Second, the constraint of relocating one prompt per iteration (see § E) restricts our method to relocating only limited prompts within certain epochs. This particularly affects tasks requiring high prompt numbers, hindering PRO-VPT’s ability to learn optimal distributions. For instance, the changes from initial uniform to learned prompt distributions are relatively less pronounced for Camelyon and DMLab as shown in Fig. ix, which require 100 initialized prompts per block. Nevertheless, it is worth noting that PRO-VPT still delivers considerable performance enhancements on these datasets. Addressing these limitations would further enhance the efficiency and applicability of our approach in diverse downstream tasks.
@@
Moreover, while this paper focuses on calibrating the distribution of prompts, the PRO-VPT framework is applicable to most block-wise PEFT approaches, e.g., adjusting the pre-block dimension of subspaces (i.e., the rank distribution) in adapter-based methods. An essential future direction deserving of further investigation is integrating our approach with other PEFT methods, developing a universal framework for fine-grained robust PEFT.